OpenStudy (anonymous):

write an equation of a line perpendicular to y=1/4x+4 in slope intercept form that passes through the point (-2,6). a) y=1/4x -4 b)y=1/4x +6 c) y=-4x -2 d) y=-4x +1/4

OpenStudy (jdoe0001):

hmm what's the slope of -> y=1/4x+4 anyway?

OpenStudy (anonymous):

1/4?

OpenStudy (jdoe0001):

yeap... so we know the slope of this equation is 1/4 so the equation of a line that is PERPENDICULAR to this one will be the NEGATIVE RECIPROCAL of that, that is \(\bf \cfrac{1}{{\color{blue}{ 4}}}\qquad negative\to - \cfrac{1}{{\color{blue}{ 4}}}\qquad reciprocal\to - \cfrac{{\color{blue}{ 4}}}{1}\to {\color{blue}{ 4}}\) so you're really asked on "what is the equation of a line with slope of "4" and that passes through (-2,6)" so \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (jdoe0001):

well... actually shold be MINUS 4 ahemm so \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (jdoe0001):

\(\bf \bf \cfrac{1}{{\color{blue}{ 4}}}\qquad negative\to - \cfrac{1}{{\color{blue}{ 4}}}\qquad reciprocal\to - \cfrac{{\color{blue}{ 4}}}{1}\to {\color{blue}{ -4}}\)

OpenStudy (anonymous):

ok so im going to plug this in

OpenStudy (jdoe0001):

yeap and then you can just solve for "y" if you need it in "slope-intercept" form

OpenStudy (anonymous):

which of the answers would it be can you solve it please so i dont get it wrong?

OpenStudy (anonymous):

because im getting a little confused

OpenStudy (jdoe0001):

well... what have you gotten so far?

OpenStudy (anonymous):

y-6=1/4(x-(-2)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -4 \\ \quad \\ y-{\color{blue}{ 6}}={\color{green}{ -4}}(x-{\color{red}{ (-2)}})\implies y-6=-4(x+2)\implies y-6=-4x-8 \\ \quad \\ y=-4x-8+6\implies y=-4x-2\)

OpenStudy (anonymous):

yay!!! thank you soooo much

OpenStudy (jdoe0001):

yw